The refractive index of glass is 1.5, if the velocity of light in air is 3×10^8 m/s, then calculate the velocity of light when it travels through glass


1
Expert's answer
2021-01-10T18:10:26-0500

By definition, refractive index is:


"n = \dfrac{c}{v}"

where "c = 3\times10^8m/s" is the speed of light in air, and "v" is the velocity of light when it travels through glass. Since "n = 1.5", obtain:


"v = \dfrac{c}{n} = \dfrac{3\times10^8m/s}{1.5} = 2\times10^8m/s"

Answer. "2\times10^8m/s".